OpenStudy (anonymous):

A pizza shop offers the topping shown below. How many three toppings pizzas can you make. pepperoni mushrooms sausage onion ham 6 10 4 5

OpenStudy (anonymous):

the number of ways to choose 3 out of 5 is \[\binom{5}{3}=\frac{5\times 4\times 3}{3\times 2}\]

OpenStudy (anonymous):

>

OpenStudy (anonymous):

6?